OpenStudy (cenaida):

If y = x + 4, simplify the expression x + 2y. A. 3x + 4 B. 3x + 8 C. 2x + 4 D. x + 4 + 2y

OpenStudy (anonymous):

x + 2(x + 4) = x + 2x + 8 = 3x + 8

OpenStudy (anonymous):

make substitution in for y = x+4 in x + 2y , then use algebra to clean up and you get your answer.

OpenStudy (anonymous):

B.